b) The volume of an object as a function of time is calculated by 

V= At3 + B/t, where t is time measured in seconds and V is in cubic meters. Determine the dimension of the constants A and B.                           

c. A vector B, when added to the vector C = 3.0i+4.0j, yields a resultant vector that is in the positive y-direction and has a magnitude equal to that of C. What is the magnitude of B? 



1
Expert's answer
2021-03-23T11:12:56-0400

b)


"[A]=L^3T^{-3}\\ [B]=L^3T"

c)


"C=\sqrt{3^2+4^2}=5\\B_x=-3\\B_y+4=5\\B_y=1\\B=\sqrt{(-3)^2+1^2}=\sqrt{10}"